OpenStudy (anonymous):

Find the exact value of sin 240°.

OpenStudy (anonymous):

sin 240 = sin (180 +60) = -sin 60 = sqrt (3)/2 = 1.732/2 = 0.8660

OpenStudy (anonymous):

\[-\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

@Bev that isnt the exact value of sin240

OpenStudy (anonymous):

Thanks, It's - sin 60 = -0.8660

OpenStudy (anonymous):

...no, the exact value is the answer i stated

OpenStudy (anonymous):

ur answer is an approximation